equation of motion of a body moving along x-axis at an instant t second is given by x = 40 + 12t  – t 3 m. displacement of the particle before coming to rest is

Grade:12th pass

1 Answers

Vara
6 years ago
Here the given equation of displacement with time is x=40+12t-t3 When we differentiate the equation we get v=12-3t2Substituting v=0 as given it comes to restWe get t=2We get x=56